OpenStudy (anonymous):

Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (abb0t):

Try squaring both sides.

OpenStudy (anonymous):

x^2+y^2=16

OpenStudy (anonymous):

you have gone from polar equations to a cartesian equation by eliminating t

OpenStudy (abb0t):

\[x^2 = 4\cos^2(t) ----- y^2=4\sin^2(t) \]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes you elimiate t to form a cartesian equation involving just x and y